Supposing a clock takes 7 seconds to strike 7 how long will the same clock take to strike 10?
my younger brother got this at school, shame on me, but can't match the answer. please help
1
Expert's answer
2014-04-23T10:08:42-0400
While striking 7 the clock stroke its first gong at 7 o'clock and it strikes 6 more at regular intervals.

These 6 intervals take 7 seconds so that the intervals between gong is 7/6 seconds. However, to strike 10 there are 9 intervals each taking 7/6 seconds, so in total we have 9*7/6=3*7/2=21/2 or 10.5 seconds.